A second year physical sciences student drives at an average speed of 40 km/h to see a friend, 

who lives a distance of 100km/h away (you may take the problem as being one-dimensional!). 

On arrival she discovers that he has misunderstood the arrangements and has already left to 

drive to visit her. She returns home, driving at an average speed of 70 km/h on the return trip. 

What is her average velocity for the whole journey? What is her average speed? If each left 

originally at the same time and his average speed is 30 km/h, who arrives first at her house? 


1
Expert's answer
2022-05-22T14:38:20-0400

"v=\frac{2v_1v_2}{v_1+v_2}=50.9~\frac{km}h."